OpenStudy (anonymous):

two plane mirrors are inclined at 70 deg. a ray incident on on one mirror at angle theta, after reflection falls on second mirror and is reflected from there parallel to the first mirror. What is theta

OpenStudy (anonymous):

Lots of geometry! See attached file. First see figure on left. (1,2,3 .... are angles, not numbers!) We know 1 is equal to θ (angle of incidence) 2=1=θ (law of reflection) 4 =70 (angle of inclination of mirrors) Now it can be proved easily (and i will prove later) that 4=8 =70 Since 7+8 =180, therefore ,7=110 Also 3+2=90 (because N1 is normal to mirror) Hence 3=90-θ Now, 3=5+6 (as second reflected ray is parallel to mirror 1) Therefore 5+6=90-θ Also 5=6 (law of reflection) Hence 5=6=45-θ/2 I have shown all the angles in the right figure. Now see triangle ABC, sum of all angle should be 180. Hence 2+5+7 =180 θ+45-θ/2+110=180 giving θ=50

OpenStudy (anonymous):

the proof that angle 4 is equal to angle 8 You can even state it as a theorem - angle between two lines is same as angle b/w their perpendiculars. N1 and N2 are perpendiculars to mirror 1 and mirror 2. So angle b/w N1 and N2 is same as the angle b/w mirror 1 and mirror 2. But which angle? There are 2 angles b/w N1 and N2 -one 70 and other 110. It may cause a confusion so it is better to find out geometrically as i have done above-I blindly applied the thm initially and wrongly persumed that angle 7=70!